OpenStudy (anonymous):

A chemist mixes a 10% hydrogen peroxide solution with a 25% hydrogen peroxide solution to create 30 liters of a 15% hydrogen peroxide solution. How many liters of the 10% solution did the chemist use to make the 15% solution?

OpenStudy (kainui):

So what does 10% of some solution mean? It means that you take the amount that's in there and divide it by the total volume of other stuff. So like if you have 10% salt solution and let's say it has 20 liters of stuff like salt and water then that means that 2 liters of the 20 liters is salt. So you write this as 2/20=10%. Try using this to help you figure this out, give it your best shot.

OpenStudy (anonymous):

I don't understand

OpenStudy (amistre64):

i would approach this with a system of equations ....

OpenStudy (amistre64):

A chemist mixes a 10% hydrogen peroxide solution with a 25% hydrogen peroxide solution to create 30 liters of a 15% hydrogen peroxide solution. let x and y represent the amounts of solution that was used such that: x+y = 30 now we can focus on the percentages used: if x is 10% and y is 25% then we know that: .10x + .25y = 30(.15) this gives us 2 equations in 2 unknowns

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

so, given: x + y = 30 .10x + .25y = 30(.15) what is your favorite method of solving a system of equations?

OpenStudy (anonymous):

Multiplication or Addition ?

OpenStudy (amistre64):

if you mean elimination .. i know they use different terms for it these days. then we can apply that. what is your idea to eliminate one of the variables?

OpenStudy (anonymous):

You subtract one of them from both sides?

OpenStudy (amistre64):

lets say we want to get rid of the x parts; if we multiply the top by say: -.10 and add the results ... -.10x -.10y = -30(.10) .10x + .25y = 30(.15) --------------------- 0x + ______________ what is the results?

OpenStudy (anonymous):

Do you multiply those?

OpenStudy (amistre64):

add (combine) like terms -.10x + .10x = 0x what do the y terms get us? the constant terms get us?

OpenStudy (anonymous):

-9.75 ?

OpenStudy (amistre64):

-.10x -.10y = -30(.10) .10x + .25y = 30(.15) --------------------- 0x + ______________ soo ... -.10x+ .10x = 0x -.10y + .25y = ?? -30(.10) + 30(.15) = ??

OpenStudy (amistre64):

if this method is to difficult, we can always try the basic substitution method

OpenStudy (anonymous):

-.10y + .25y = 0.15 -30 * .10 = -3 30 * .15 = 4.5

OpenStudy (anonymous):

-3 + 4.5 = 1.5 ?

OpenStudy (amistre64):

correct so far :) -.10x -.10y = -30(.10) .10x + .25y = 30(.15) --------------------- .15y = 1.5 now we can solve for y

OpenStudy (anonymous):

1.5 divided by .15 = 10

OpenStudy (amistre64):

so, y=10 and since x + y = 30, what is x?

OpenStudy (anonymous):

20 ?

OpenStudy (amistre64):

yeah so we now know what the solution is x + y = 30 20+10 = 30 .10x + .25y = 30(.15) 20L of 10% and 10L of 25%

OpenStudy (anonymous):

So the answer is 20 liters ?

OpenStudy (amistre64):

How many liters of the 10% solution did the chemist use to make the 15% solution? we need both results

OpenStudy (anonymous):

Yeah , its 20. Thank you so much

OpenStudy (amistre64):

youre welcome :) good luck

OpenStudy (amistre64):

just for the record ... @igreen any thoughts?

OpenStudy (igreen):

Looks great!